# PEneo: Unifying Line Extraction, Line Grouping, and Entity Linking for End-to-end Document Pair Extraction

### v1

@@ Abstract
Document pair extraction aims to identify key and value entities as well as their relationships from visually-rich documents. Most existing methods divide it into two separate tasks: semantic entity recognition (SER) and relation extraction (RE). However, simply concatenating SER and RE serially can lead to severe error propagation, and it fails to handle cases like multi-line entities in real scenarios. To address these issues, this paper introduces a novel framework, PEneo (Pair Extraction new decoder option), which performs document pair extraction in a unified pipeline, incorporating three concurrent sub-tasks: line extraction, line grouping, and entity linking. This approach alleviates the error accumulation problem and can handle the case of multi-line entities. Furthermore, to better evaluate the model’s performance and to facilitate future research on pair extraction, we introduce RFUND, a re-annotated version of the commonly used FUNSD and XFUND datasets, to make them more accurate and cover realistic situations. Experiments on various benchmarks demonstrate PEneo’s superiority over previous pipelines, boosting the performance by a large margin (e.g., 19.89%-22.91% F1 score on RFUND-EN) when combined with various backbones like LiLT and LayoutLMv3, showing its effectiveness and generality. Codes and the new annotations will be open to the public.
1 Introduction
@@
Document pair extraction is a vital step in analyzing form-like documents containing information organized as key-value pairs. It involves identifying the key and value entities, as well as their linking relationships from document images. Previous research has generally divided it into two document understanding tasks: semantic entity recognition (SER) and relation extraction (RE). The SER task involves extracting contents that belong to predefined categories, such as retrieving store names and prices from receipts [11] or analyzing nutrition facts labels [14]. Most of the existing methods [26, 28, 27, 10, 23, 16] implement SER using BIO tagging, where tokens in the input text sequence are tagged as the beginning (B), inside (I), or outside (O) element for each entity. On the other hand, the RE task aims to identify relations between given entities, such as predicting the linkings between form elements [7, 29]. Previous works [28, 27, 23] have typically employed a linking classification network for relation extraction: given the entities in the document, it first generates representations for all possible entity pairs, then applies binary classification to filter out the valid ones. Document pair extraction is usually achieved by serially concatenating the above two tasks (SER+RE), where the SER model first identifies all the key and value entities from the document, and the RE model finds the matching values for each key. Figure 1 provides examples of SER, RE, and document pair extraction.
@@
Although achievements have been made in SER and RE, the existing SER+RE approach overlooks several issues. In previous settings, SER and RE are viewed as two distinct tasks that have inconsistent input/output forms and employ simplified evaluation metrics. For the SER task, entity-level OCR results are usually given [7], where text lines belonging to the same entity are aggregated and serialized in human reading order. The model categorizes each token based on the well-organized sequence, neglecting the impact of improper OCR outputs. In the RE task, models take the ground truths of the SER task as input, using prior knowledge of entity content and category. The model simply needs to predict the linkings based on the provided key and value entities, and the linking-level F1 score is taken as the evaluation metric. In real-world applications, however, the situation is considerably more complex. Commonly used OCR engines typically generate results at the line level. For entities with multiple lines, an extra line grouping step is required before BIO tagging, which is hard to realize for complex layout documents. Additionally, errors in SER predictions can significantly impact the RE step, resulting in unsatisfactory pair extraction results. Section 5.5 analyzes the SER+RE performance drop in detail.
@@
Furthermore, we found that some annotations in the two commonly used form understanding datasets, FUNSD [7] and XFUND [29], do not meet the real-world requirements well. Hence we propose their relabeled version, RFUND. The inconsistent labeling granularity in the original annotations is unified into line-level, aiming to imitate the output of a real OCR engine. We also rectified the category and relation annotations to make it more clear.
@@
2 Related Work
@@
Early studies [25, 22] utilized heuristic rules to extract key-value pairs from documents. These approaches exhibit limited applicability to specific document layouts and demonstrate poor generalization performance. In recent years, with the advancements in deep learning techniques, researchers have proposed several deep learning-based methods for pair extraction. LayoutLM [26] has first proposed embedding the coordinate of each word into a BERT-style model to capture the multi-modal features of each token. It also strengthens the model’s representational capability with specially designed pre-training tasks. Subsequent works primarily focus on improving the backbones to obtain more powerful and general token representations. [8] introduces a novel relative spatial encoding to capture layout information effectively. [28, 10, 16, 15, 1, 5] incorporate visual features and enhance the interaction of different modalities through newly designed architectures and pre-training tasks. [23] proposes a two-branch structure that allows flexible switching of semantic encoding modules and fast adaptation in different language scenarios. To handle the text serialization problem, ERNIE-Layout [21] employs a Layout-Parser as well as a reading order prediction task to recognize and sort the text segments. To boost performance on RE, GeoLayoutLM [19] proposes a novel relation head and a geometric pre-training schema, and obtains outstanding performance on various RE benchmarks [29, 20]. ESP [30] introduces an end-to-end pipeline incorporating text detection, text recognition, entity extraction, and entity linking. It also predicts inter/intra-linkings between words to cope with multi-line entities. The model shows outstanding performance on various relation extraction tasks. However, all of the aforementioned methods achieve pair extraction by simply concatenating the downstream SER and RE model, thereby overlooking the error accumulations in this process.
@@
In addition to the SER+RE pipeline, some other approaches explore alternative ways to accomplish pair extraction. FUDGE [3] employs a graph-based scheme that iteratively aggregates text lines and predicts the key-value linking. Although it is capable of handling multi-line entities, its performance is relatively limited due to the absence of a powerful multi-modal encoder. SPADE [12] handles the document parsing task using the dependency parsing strategy, and it demonstrates good performance on the CORD [20] dataset. However, it models and decodes the document at the word level and thus has a huge computational overhead. Donut [13] and Dessurt [4] both propose image-to-sequence pipelines, which achieve pair extraction in a question-answering manner; QGN [2] introduces a query-driven network that generates the pair predictions using the value prefix. These generative models require a lot of training data and fail to handle complex layout documents. DocTr [17] first identifies anchor words from the input OCR results, then predicts entity-level bounding boxes and relations using a vision-language decoder, achieving structured information extraction and multi-line entity grouping. However, it requires task-specific pre-training and cannot be directly applied to existing backbones. KVPFormer [9] takes a different approach by first identifying key entities in the given documents and then employing an answer prediction module to determine their corresponding values. Notably, KVPFormer achieves exceptional performance in the RE task without pre-training, but it requires prior knowledge of entity spans and cannot handle unordered OCR inputs.
3 The RFUND Dataset
3.1 Analysis of the Original Annotations
@@
The FUNSD dataset is a commonly used form understanding benchmark that comprises scanned English documents. The XFUND dataset is its multilingual extension, covering 7 languages (Chinese, Japanese, Spanish, French, Italian, German, and Portuguese). Entities in these forms are categorized into four types, including header, question, answer, and other. Linking relationships between different entities are annotated to represent the structure of the form.
@@
While most contents in FUNSD are annotated at the entity level, multi-line entities with first-line indentation are annotated separately, as illustrated in Figure 1(a). XFUND exhibits variable granularity in annotations, with some contents labeled at the entity level and others at the line level, as shown in Figure 1(b). Such inconsistent labeling standards can hinder model training and fail to work in real-world scenarios. Moreover, it is observed that certain entities in both FUNSD and XFUND have category labels that differ from human understanding (Figure 1(c)), highlighting the need for refinement of the annotations in this aspect.
@@
3.2 Relabeling for Real-World Scenarios
Given the original entity-level annotations, we employed rules to generate line-level text strings and bounding boxes. For cases that could not be handled perfectly, we made manual corrections. We then grouped text lines belonging to the same entity, including the cases of first-line indent described above. Any redundant linking labels were removed. We also rectified the entity category labels and key-value linkings to make them consistent with human understanding. We also eliminated header-to-question linkings that describe nested information, simplifying the task scope. We term the resulting dataset as RFUND, and Table 1 summarizes its key statistics.
3.3 Task Definition
The model is expected to take the line-level OCR results, which are commonly used in real scenarios, as input. It should then predict all the key-value pairs in string format. Pair-level F1 score is employed for performance comparison, where a pair prediction will be considered as True Positive if and only if both the predicted key and value text strings match the ground truths exactly.
4 PEneo
@@
4.1 Multi-modal Encoder
The encoder tokenizes the input text lines into tokens and integrates semantic, layout, and visual (optional) information to obtain multi-modal features for each token. Various BERT-like document understanding models, including LayoutLMv2 [28], LayoutLMv3 [10], and LiLT [23], can serve as the encoding backbone for PEneo. To reduce memory consumption for the following operation, we append a linear projection layer to map the channels of the output features to a smaller size at the final stage:
@@ 4.1 Multi-modal Encoder
4.2 Joint Extraction Decoder
@@
Inspired by [24], token representations from the encoder are concatenated in a pair-wise manner. Subsequently, a pair encoding layer is applied to obtain the token pair representations matrix , where is the number of input tokens. Each entry is computed as
@@ 4.2.3 Entity Linking
4.2.4 Linking Parsing
Matrix predicts the first token of each key and value entity. The Span of the entity’s first line can be determined by referring to the line extraction result from . For multi-line entities, based on the first and last token of each line, spans of the entity’s following lines can be retrieved iteratively from the line grouping predictions and . Once we collect all the contents of the current pair, we compare the last token of the key and value entity with the entity tail linking result to determine the validity of the current prediction. During the parsing process, if the predictions of different matrices are found to be contradictory, the current parsed content is considered to be erroneous and is directly eliminated. The algorithm flow of the linking parsing module is discussed in the supplementary materials.
4.3 Supervised Learning Target
@@ 5.1 Datasets
We conduct experiments on RFUND and SIBR [30]. As described in section 3, RFUND contains 8 subsets corresponding to 8 different languages. We follow the language-specific fine-tuning settings in [27] to evaluate the model’s performance on each subset. SIBR is a bilingual dataset composed of 600 training samples and 400 testing samples. It contains 600 Chinese invoices, 300 English bills of entry, and 100 bilingual receipts. The dataset is annotated at line level, with entity linking (inter-links) and line grouping (intra-links) labels provided. We observed some contradictory annotations in SIBR and made manual corrections. In our experiment, we focus on the linkings between question and answer entities in SIBR and employ pair-level F1 score as the evaluation metric.
5.2 Implementation Details
@@
The reduced feature channel size is set to . During the training phase, the loss weighting hyper-parameters are all set to 1. To address category imbalance, the class weighting tensor for the cross-entropy loss is set to . We employ AdamW [18] as the optimizer. The learning rate is set to 2e-6 for the encoder backbone and 1e-4 for the decoder, scheduled by a linear scheduler with a warm-up ratio of 0.1. We fine-tune PEneo for 650 epochs on RFUND and 330 epochs on SIBR, with a batch size of 4.
5.3 Baseline Settings
@@
We employed several widely used and publicly available models, LiLT, LayoutLMv2, LayoutXLM, and LayoutLMv3, as the encoding backbone to evaluate the effectiveness of our proposed PEneo framework. The baseline method serially combines an SER model and a RE model for pair extraction. At the SER stage, we sorted the input lines with Augmented XY Cut [6], aiming to maximize the adjacency of lines within an entity. The model learns to extract entities based on the sorted sequence and group the correctly ordered lines through BIO tagging. For the RE part, we train the model using entity-level annotations, consistent with previous studies. Additionally, we considered GeoLayoutLM [19], which includes a powerful RE decoder, as the baseline. We perform pair extraction by concatenating its downstream SER and RE models, following the aforementioned settings. Furthermore, we adopt Donut [13], an OCR-free end-to-end model with an image-to-sequence pipeline, to our pair extraction task. It takes the document image as input and predicts HTML-like strings representing the document structure. It is worth noting that some models only provide pre-trained weights for a single or a small number of languages, which does not cover all the samples in RFUND. Hence, we only evaluated the model’s performance on the language subset covered by their pre-training corpus. Further details are discussed in supplementary materials.
Method | Pipeline | F1 |
Donut [13] | End-to-End | 24.54 |
GeoLayoutLM [19] | SER+RE | 69.03 |
LiLT[EN-R] [23] | SER+RE | 54.33 |
PEneo-LiLT[EN-R] | Joint | 74.22 (+19.89) |
LiLT[InfoXLM] [23] | SER+RE | 52.18 |
PEneo-LiLT[InfoXLM] | Joint | 74.29 (+22.11) |
LayoutXLM [27] | SER+RE | 52.98 |
PEneo-LayoutXLM | Joint | 74.25 (+21.27) |
LayoutLMv2 [28] | SER+RE | 49.06 |
PEneo-LayoutLMv2 | Joint | 71.97 (+22.91) |
LayoutLMv3 [10] | SER+RE | 57.66 |
PEneo-LayoutLMv3 | Joint | 79.27 (+21.61) |
Method | Pipeline | F1 |
Donut [23] | End-to-End | 17.26 |
LiLT[InfoXLM] [23] | SER+RE | 72.76 |
PEneo-LiLT[InfoXLM] | Joint | 82.36 (+9.60) |
LayoutXLM [27] | SER+RE | 70.45 |
PEneo-LayoutXLM | Joint | 82.23 (+11.78) |
LayoutLMv3 [10] | SER+RE | 73.51 |
PEneo-LayoutLMv3 | Joint | 82.52 (+9.01) |

### v2

@@ Abstract.
Document pair extraction aims to identify key and value entities as well as their relationships from visually-rich documents. Most existing methods divide it into two separate tasks: semantic entity recognition (SER) and relation extraction (RE). However, simply concatenating SER and RE serially can lead to severe error propagation, and it fails to handle cases like multi-line entities in real scenarios. To address these issues, this paper introduces a novel framework, PEneo (Pair Extraction new decoder option), which performs document pair extraction in a unified pipeline, incorporating three concurrent sub-tasks: line extraction, line grouping, and entity linking. This approach alleviates the error accumulation problem and can handle the case of multi-line entities. Furthermore, to better evaluate the model’s performance and to facilitate future research on pair extraction, we introduce RFUND, a re-annotated version of the commonly used FUNSD and XFUND datasets, to make them more accurate and cover realistic situations. Experiments on various benchmarks demonstrate PEneo’s superiority over previous pipelines, boosting the performance by a large margin (e.g., 19.89%-22.91% F1 score on RFUND-EN) when combined with various backbones like LiLT and LayoutLMv3, showing its effectiveness and generality. Codes and the new annotations are available at https://github.com/ZeningLin/PEneo.
1. Introduction
@@
Document pair extraction is a vital step in analyzing form-like documents containing information organized as key-value pairs. It involves identifying the key and value entities, as well as their linking relationships from document images. Previous research has generally divided it into two document understanding tasks: semantic entity recognition (SER) and relation extraction (RE). The SER task involves extracting contents that belong to predefined categories, such as retrieving store names and prices from receipts (Huang et al., 2019) or analyzing nutrition facts labels (Kuang et al., 2023). Most of the existing methods (Xu et al., 2020, 2021b, 2021a; Huang et al., 2022; Wang et al., 2022; Li et al., 2021b) implement SER using BIO tagging, where tokens in the input text sequence are tagged as the beginning (B), inside (I), or outside (O) element for each entity. On the other hand, the RE task aims to identify relations between given entities, such as predicting the linkings between form elements (Guillaume Jaume, 2019; Xu et al., 2022). Previous works (Xu et al., 2021b, a; Wang et al., 2022) have typically employed a linking classification network for relation extraction: given the entities in the document, it first generates representations for all possible entity pairs, then applies binary classification to filter out the valid ones. Document pair extraction is usually achieved by serially concatenating the above two tasks (SER+RE), where the SER model first identifies all the key and value entities from the document, and the RE model finds the matching values for each key. Figure 1 provides examples of SER, RE, and document pair extraction.
@@
Although achievements have been made in SER and RE, the existing SER+RE approach overlooks several issues. In previous settings, SER and RE are viewed as two distinct tasks that have inconsistent input/output forms and employ simplified evaluation metrics. For the SER task, entity-level OCR results are usually given (Guillaume Jaume, 2019), where text lines belonging to the same entity are aggregated and serialized in human reading order. The model categorizes each token based on the well-organized sequence, neglecting the impact of improper OCR outputs. In the RE task, models take the ground truths of the SER task as input, using prior knowledge of entity content and category. The model simply needs to predict the linkings based on the provided key and value entities, and the linking-level F1 score is taken as the evaluation metric. In real-world applications, however, the situation is considerably more complex. Commonly used OCR engines typically generate results at the line level. For entities with multiple lines, an extra line grouping step is required before BIO tagging, which is hard to realize for complex layout documents. Additionally, errors in SER predictions can significantly impact the RE step, resulting in unsatisfactory pair extraction results. Section 5.5 analyzes the SER+RE performance drop in detail.
@@
Furthermore, we found that some annotations in the two commonly used form understanding datasets, FUNSD (Guillaume Jaume, 2019) and XFUND (Xu et al., 2022), do not meet the real-world requirements well. Hence we propose their relabeled version, RFUND. The inconsistent labeling granularity in the original annotations is unified into line-level, aiming to imitate the output of a real OCR engine. We also rectified the category and relation annotations to make it more clear.
@@
2. Related Work
2.1. Document Pair Extraction Methods
@@
Early studies (Watanabe et al., 1995; Seki et al., 2007) utilized heuristic rules to extract key-value pairs from documents. These approaches exhibit limited applicability to specific document layouts and demonstrate poor generalization performance. In recent years, with the advancements in deep learning techniques, researchers have proposed several deep learning-based methods for pair extraction. LayoutLM (Xu et al., 2020) has first proposed embedding the coordinate of each word into a BERT-style model to capture the multi-modal features of each token. It also strengthens the model’s representational capability with specially designed pre-training tasks. Subsequent works primarily focus on improving the backbones to obtain more powerful and general token representations. (Hong et al., 2022) introduces a novel relative spatial encoding to capture layout information effectively. (Xu et al., 2021b; Huang et al., 2022; Li et al., 2021b, a; Appalaraju et al., 2021; Gu et al., 2021) incorporate visual features and enhance the interaction of different modalities through newly designed architectures and pre-training tasks. (Wang et al., 2022) proposes a two-branch structure that allows flexible switching of semantic encoding modules and fast adaptation in different language scenarios. To handle the text serialization problem, ERNIE-Layout (Peng et al., 2022) employs a Layout-Parser as well as a reading order prediction task to recognize and sort the text segments. To boost performance on RE, GeoLayoutLM (Luo et al., 2023) proposes a novel relation head and a geometric pre-training schema, and obtains outstanding performance on various RE benchmarks (Xu et al., 2022; Park et al., 2019). ESP (Yang et al., 2023) introduces an end-to-end pipeline incorporating text detection, text recognition, entity extraction, and entity linking. It also predicts inter/intra-linkings between words to cope with multi-line entities. The model shows outstanding performance on various relation extraction tasks. However, all of the aforementioned methods achieve pair extraction by simply concatenating the downstream SER and RE model, thereby overlooking the error accumulations in this process.
@@
In addition to the SER+RE pipeline, some other approaches explore alternative ways to accomplish pair extraction. FUDGE (Davis et al., 2021) employs a graph-based detection scheme that iteratively aggregates text lines and predicts the key-value linking. Although it is capable of handling multi-line entities, its performance is relatively limited due to the absence of semantic information. SPADE (Hwang et al., 2021) handles the document parsing task using the dependency parsing strategy, and it demonstrates good performance on the CORD (Park et al., 2019) dataset. However, it models and decodes the document based on quantities of word-to-word relation, resulting in a huge computational overhead. Donut (Kim et al., 2022) and Dessurt (Davis et al., 2022) both propose image-to-sequence pipelines, which achieve pair extraction in a question-answering manner; QGN (Cao et al., 2022) introduces a query-driven network that generates the pair predictions using the value prefix. These generative models require a lot of training data and fail to handle complex layout documents. DocTr (Liao et al., 2023) first identifies anchor words from the input OCR results, then predicts entity-level bounding boxes and relations using a vision-language decoder, achieving structured information extraction and multi-line entity grouping. However, it requires task-specific pre-training and cannot be directly applied to existing backbones. TPP (Zhang et al., 2023) employs a unified token path prediction framework for multi-line SER and RE tasks, and it outperforms conventional BIO-tagging baselines. However, TPP regards RE as a token clustering task, which leads to the inability to differentiate the key and value content individually. KVPFormer (Hu et al., 2023) takes a different approach by first identifying key entities in the given documents and then employing an answer prediction module to determine their corresponding values. Notably, its proposed spatial compatibility feature helps achieve exceptional performance in the RE task without pre-training, but it requires prior knowledge of entity spans and cannot handle unordered OCR inputs.
2.2. Joint Extraction in Plain Texts
Joint extraction aims to identify the subject-relation-object triplets simultaneously from plain texts. Mainstream approaches can be broadly categorized into two types: pipeline-based methods and joint methods.
The pipeline-based method is akin to the SER+RE approach mentioned above. It involves a sequential combination of SER and RE tasks, wherein the entity contents are initially predicted, followed by the classification of the semantic relation type between them (Wang et al., 2019; Soares et al., 2019; Zhong and Chen, 2021; Ye et al., 2022). The joint methods unify the two tasks through specifically designed network architecture. (Zheng et al., 2017; Dai et al., 2019) propose novel tagging schemes that incorporate relation extraction annotations into BIO tags, thereby unifying the entire pipeline in BIO tagging manners. (Wei et al., 2020) employs span prediction to extract the subject content from the token sequence, subsequently predicting their corresponding objects and relation types through relation-specific object taggers. (Wang et al., 2020) leverages span prediction for entity identification, then constructs head and tail linking matrices for relation extraction. This approach enables the parsing of triplets through a joint decoding schema.
It is noteworthy that while there are valuable insights to be gained from joint extraction in plain texts, the extraction of key-value pairs in visually-rich documents presents unique challenges that necessitate additional efforts. For instance, the entities to be extracted may span across multiple OCR boxes, giving rise to challenges related to multi-line SER. Moreover, the determination of relationships between entities relies on spatial information, thus requiring the development of specialized modules to effectively address these requirements.
3. The RFUND Dataset
3.1. Analysis of the Original Annotations
@@
The FUNSD dataset is a commonly used form understanding benchmark that comprises scanned English documents. The XFUND dataset is its multilingual extension, covering 7 languages (Chinese, Japanese, Spanish, French, Italian, German, and Portuguese). Entities in these forms are categorized into four types, including header, question, answer, and other. Entity-level and word-level OCR results are provided, and linking relationships between different entities are annotated to represent the structure of the form.
@@
While most contents in FUNSD are annotated at the entity level, multi-line entities with first-line indentation are annotated in a distinct manner. As illustrated in Figure 2(a), the first line of the answer paragraph is considered a separate entity, while the other lines remain aggregated as another. The question entity is linked to both answers, leading to redundant annotations. XFUND exhibits variable granularity in annotations, with some contents labeled at the entity level and others at the line level, as shown in Figure 2(b). Such inconsistent labeling standards can hinder model training and fail to work in real-world scenarios. Moreover, it is observed that certain entities in both FUNSD and XFUND have category labels that differ from human understanding (Figure 2(c)), highlighting the need for refinement of the annotations in this aspect.
3.2. Relabeling for Real-World Scenarios
Based on the original entity-level and word-level OCR results, we implemented a set of rules to divide paragraphs into line-level text strings and bounding boxes. Specifically, we examined the vertical distance between two adjacent words within an entity. If the difference exceeds the average height of entity words, we assign the latter word to the subsequent line. For cases that could not be handled perfectly by the rules, we made manual corrections. To represent entity-level information, line grouping annotations were added to indicate the correct order of line aggregation. Cases of first-line indent described above were corrected, and any redundant linking labels were removed. To ensure consistency with human understanding, we adjusted the entity category labels and key-value linkings accordingly. In addition, we eliminated header-to-question linkings that describe nested information, simplifying the task scope. We term the resulting dataset as RFUND, and Table 1 summarizes its key statistics.
3.3. Task Definition
The model is expected to take the line-level OCR results, which are commonly used in real scenarios, as input. It should then predict all the key-value pairs in string format. Pair-level F1 score is employed for performance comparison, where a pair prediction will be considered as True Positive if and only if both the predicted key and value text strings exactly match the ground truths.
@@ 3.3. Task Definition
4. PEneo
@@
4.1. Multi-modal Encoder
The encoder tokenizes the input text lines into tokens and integrates semantic, layout, and visual (optional) information to obtain multi-modal features for each token. Various BERT-like document understanding models, including LayoutLMv2 (Xu et al., 2021b), LayoutLMv3 (Huang et al., 2022), and LiLT (Wang et al., 2022), can serve as the encoding backbone for PEneo. To reduce memory consumption for the following operation, we append a linear projection layer to map the channels of the output features to a smaller size at the final stage:
@@ 4.1. Multi-modal Encoder
4.2. Joint Extraction Decoder
@@
Inspired by (Wang et al., 2020), token representations from the encoder are concatenated in a pair-wise manner. Subsequently, a pair encoding layer is applied to obtain the token pair representations matrix , where is the number of input tokens. Each entry is computed as
@@ 4.2.3. Entity Linking
4.2.4. Linking Parsing
Matrix predicts the first token of each key and value entity. The Span of the entity’s first line can be determined by referring to the line extraction result from . For multi-line entities, based on the first and last token of each line, spans of the entity’s following lines can be retrieved iteratively from the line grouping predictions and . Once we collect all the contents of the current pair, we compare the last token of the key and value entity with the entity tail linking result to determine the validity of the current prediction. During the parsing process, if the predictions of different matrices are found to be contradictory, the current parsed content is considered to be erroneous and is directly eliminated.
4.3. Supervised Learning Target
@@ 5.1. Datasets
We conduct experiments on RFUND and SIBR (Yang et al., 2023). As described in section 3, RFUND contains 8 subsets corresponding to 8 different languages. We follow the language-specific fine-tuning settings in (Xu et al., 2021a) to evaluate the model’s performance on each subset. SIBR is a bilingual dataset composed of 600 training samples and 400 testing samples. It contains 600 Chinese invoices, 300 English bills of entry, and 100 bilingual receipts. The dataset is annotated at line level, with entity linking (inter-links) and line grouping (intra-links) labels provided. We observed some contradictory annotations in SIBR and made manual corrections. In our experiment, we focus on the linkings between question and answer entities in SIBR and employ pair-level F1 score as the evaluation metric.
5.2. Implementation Details
@@
The reduced feature channel size is set to . During the training phase, the loss weighting hyper-parameters are all set to 1. To address category imbalance, the class weighting tensor for the cross-entropy loss is set to . We employ AdamW (Loshchilov and Hutter, 2019) as the optimizer. The learning rate is set to 2e-6 for the encoder backbone and 1e-4 for the decoder, scheduled by a linear scheduler with a warm-up ratio of 0.1. We fine-tune PEneo for 650 epochs on RFUND and 330 epochs on SIBR, with a batch size of 4.
5.3. Baseline Settings
@@
We employed several widely used and publicly available models, LiLT, LayoutLMv2, LayoutXLM, and LayoutLMv3, as the encoding backbone to evaluate the effectiveness of our proposed PEneo framework. The baseline method serially combines an SER model and a RE model for pair extraction. At the SER stage, we sorted the input lines with Augmented XY Cut (Gu et al., 2022), aiming to maximize the adjacency of lines within an entity. The model learns to extract entities based on the sorted sequence and group the correctly ordered lines through BIO tagging(Ramshaw and Marcus, 1999). For the RE part, we train the model using entity-level annotations, consistent with previous studies. Additionally, we adapted FUDGE (Davis et al., 2021), Donut (Kim et al., 2022), GeoLayoutLM (Luo et al., 2023), TPP (Zhang et al., 2023) and GPT-4V (Achiam et al., 2023) to the pair extraction task. FUDGE employs a graph-based detection pipeline that predicts the key-value bounding box pairs. Since it only predicts boxes, we report its performance on the box-pair F1-score as a compromise. Donut takes the document image as input and predicts HTML-like strings containing key-value pairs. GeoLayoutLM is a strong baseline that includes a powerful RE decoder, and we perform pair extraction by concatenating its downstream SER and RE models, following the aforementioned SER+RE settings. TPP employs a token clustering scheme, and the groups of pair tokens can be obtained by parsing its VrD-EL matrix with depth-first searching. Its performance is reported on token-group F1-score. For GPT-4V, we follow the evaluation pipeline proposed by (Shi et al., 2023). It is worth noting that some models only provide pre-trained weights for a single or a small number of languages, which does not cover all the samples in RFUND. Hence, we only evaluated the model’s performance on the language subset covered by their pre-training corpus.
Method | Venue | Pipeline | F1 |
FUDGE (Davis et al., 2021) | ICDAR’21 | End-to-End | 53.15‡ |
Donut (Kim et al., 2022) | ECCV’22 | Image2Seq | 24.54 |
GeoLayoutLM (Luo et al., 2023) | CVPR’23 | SER+RE | 69.03 |
TPP-LayoutLMv3BASE (Zhang et al., 2023) | EMNLP’23 | Joint | 50.27‡ |
GPT-4V w/o OCR (Achiam et al., 2023) | arXiv’23 | Image2Seq | 20.96 |
GPT-4V w OCR (Achiam et al., 2023) | arXiv’23 | Image2Seq | 38.15 |
LiLT[EN-R]BASE (Wang et al., 2022) | ACL’22 | SER+RE | 54.33 |
PEneo-LiLT[EN-R]BASE | Ours | Joint | 74.22 (+19.89) |
LiLT[InfoXLM]BASE (Wang et al., 2022) | ACL’22 | SER+RE | 52.18 |
PEneo-LiLT[InfoXLM]BASE | Ours | Joint | 74.29 (+22.11) |
LayoutXLMBASE (Xu et al., 2021a) | arXiv’21 | SER+RE | 52.98 |
PEneo-LayoutXLMBASE | Ours | Joint | 74.25 (+21.27) |
LayoutLMv2BASE (Xu et al., 2021b) | ACL’21 | SER+RE | 49.06 |
PEneo-LayoutLMv2BASE | Ours | Joint | 71.97 (+22.91) |
LayoutLMv3BASE (Huang et al., 2022) | ACMMM’22 | SER+RE | 57.66† |
PEneo-LayoutLMv3BASE | Ours | Joint | 79.27 (+21.61) |
Method | Venue | Pipeline | F1 |
DonutBASE (Wang et al., 2022) | ECCV’22 | Image2Seq | 17.26 |
LiLT[InfoXLM]BASE (Wang et al., 2022) | ACL’22 | SER+RE | 72.76 |
PEneo-LiLT[InfoXLM]BASE | Ours | Joint | 82.36 (+9.60) |
LayoutXLMBASE (Xu et al., 2021a) | arXiv’21 | SER+RE | 70.45 |
PEneo-LayoutXLMBASE | Ours | Joint | 82.23 (+11.78) |
LayoutLMv3 (Huang et al., 2022) | ACMMM’22 | SER+RE | 73.51† |
PEneo-LayoutLMv3 | Ours | Joint | 82.52 (+9.01) |